# General Hamiltonian description of nonreciprocal interactions

### v3

@@ C.1 Langevin steady-state distribution
with conjugate degrees of freedom $\{\theta_{i},\varphi_{j}\}{=}\delta_{ij}$. To maintain consistency with the traditional form of Langevin dynamics and represent the equations of motion on the constraint surface, the notation $\partial_{\theta_{i}}H|_{\text{constr.}}$ is introduced to replace the Poisson bracket. Here $\eta_{i}(t)$ denotes white noise with $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$, with $T$ the temperature of the bath, where the average is taken over the phase-space density $\rho(\theta_{i}(t),\varphi_{i}(t))$. Note that the constraint implies that each pair $(\theta_{i},\varphi_{i})$ experiences the same noise realization.
Below, we first consider steady-state distributions (i.e., no explicit $t$-dependence in $\rho$, Sec. C.1.1), and then generalize the result to non-stationary distributions in Sec. C.1.2.
C.1.1 Steady-state distribution
To obtain the steady-state phase-space density distribution $\rho(\theta_{i}(t),\varphi_{i}(t))$ under the above dynamics, we follow the standard procedure [54, 55] and introduce a test function $f(\theta_{i}(t),\varphi_{i}(t))$ over phase space. The change of the test function within a small time window $\Delta t$ is to leading order
$f\left(\theta_{i}(t+\Delta t),{\varphi}_{i}(t+\Delta t)\right)-f\left(\theta_{i}(t),{\varphi}_{i}(t)\right)$ (74)
@@ C.1.1 Steady-state distribution
$\int\frac{\mathrm{d}f\left(\theta_{i},\varphi_{i}\right)}{\mathrm{d}t}\rho(\theta_{i},\varphi_{i})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=\lim_{\Delta t\to 0}\int\frac{f\left(\theta_{i}(t+\Delta t),{\varphi}_{i}(t+\Delta t)\right)-f\left(\theta_{i}(t),{\varphi}_{i}(t)\right)}{\Delta t}\rho(\theta_{i},\varphi_{i})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=0.$ (76)
To proceed with the derivation, we write the noise terms in Eq. (C.1) as a Wiener process $W_{i}(t)$:
@@
$\int\frac{\partial f}{\partial\theta_{i}}\dot{\theta_{i}}\Delta t\;\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}\mathbf{\theta}_{i}\mathrm{d}\mathbf{\varphi}_{i}$ $=$ $\int\frac{\partial f}{\partial\theta_{i}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$ (78)
$=$ $\Delta t\int\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}},$
@@
$\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\dot{\theta_{i}}\dot{\theta_{j}}\Delta t^{2}\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}=$ (79)
$=$ $\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\left[\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{j}(t)\right]\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\int 2T\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\Delta W_{i}(t)\Delta W_{j}(t)\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\Delta W_{i}(t)\rho\left(\theta_{i},\varphi_{i}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta W_{j}(t)\rho\left(\theta_{i},\varphi_{i}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}+\Delta t^{2}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\right)\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\Delta t\delta_{ij}2T\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\rho\left(\theta_{i},\varphi_{i},t\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+O\left[\left(\Delta t\right)^{2}\right],$
@@
$+T\left(\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{i}}+\frac{\partial^{2}f}{\partial\varphi_{i}\partial\varphi_{i}}+2\frac{\partial^{2}f}{\partial\varphi_{i}\partial\theta_{i}}\right)\Bigg{]}\rho\left(\theta_{i},\varphi_{i}\right)$ $=$ $0\,.$ (80)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[\frac{\partial f}{\partial\theta_{i}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\right)-\frac{\partial f}{\partial\varphi_{i}}\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\theta_{i},\varphi_{i}\right)=0.$ (81)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[-\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)f\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\theta_{i},\varphi_{i}\right)=0.$ (82)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}f\left[\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho\left(\theta_{i},\varphi_{i}\right)\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}\rho\left(\theta_{i},\varphi_{i}\right)\right]=0.$ (83)
@@
$\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left[\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\ \rho\left(\theta_{i},\varphi_{i}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\rho\left(\theta_{i},\varphi_{i}\right)\right]=0.$ (84)
So far the derivation was generic since we did not make use of the precise form of the equations of motion. We will show in Sec. C.2 that Eq. (84) governs also the constrained Glauber dynamics.
C.1.2 Time-dependent non-stationary distribution
Before we do that, let us complete the analysis and examine the case where the density distribution exhibits explicit time dependence, $\rho(\theta_{i}(t),\varphi_{i}(t),t)$. Such situations can occur away from equilibrium where oscillating states arise.
@@ C.1.2 Time-dependent non-stationary distribution
Setting Eq.(C.1.2) equal to Eq.(87) yields
@@
This Fokker-Planck equation governs the dynamics of the phase-space distribution. As expected, for a stationary distribution, $\partial_{t}\rho=0$, it reduces to Eq. (84).
C.2 Glauber dynamics description
We now present two equivalent Glauber algorithms that sample the steady state of the Langevin dynamics. In particular, Sec. C.2.1 presents the algorithm used in the main text, while Sec. C.2.2 presents an alternative, yet equivalent, Glauber algorithm based on global balance.
C.2.1 Individual-update Glauber scheme: Detailed balance followed by imposing the constraint
In this subsection, we introduce our constrained Glauber dynamics, which is based on the Hamiltonian embedding that we propose for nonreciprocal systems. It is a Markov process with the constraint that every state $(\theta_{i},\varphi_{i})$ satisfies the condition $\theta_{i}=\varphi_{i}+\pi$. Similar to the Langevin description in the previous section, we assume that the dynamics eventually reaches a steady state, with the density distribution given by $\rho(\theta,\varphi)$.
Note that, evaluating the Hamiltonian on the constraint, we find $H(\theta_{i},\varphi_{i})=0$, which makes it difficult to define a proper Glauber update rule. Therefore, the algorithm starts by first proposing an update of only the original spin, changing its angle as $\theta\rightarrow\theta^{\prime}=\theta+\delta$. The transition rates $w(\theta\rightarrow\theta^{\prime})$ of this proposed update are calculated based on the energy difference between the original and updated state, obtained from the Hamiltonian. Subsequently, the auxiliary spin is rotated accordingly as $\varphi\rightarrow\varphi+\delta$ to enforce the mirror constraint. In practice, enforcing the constraint only in the second step allows us to define transition rates using the energy fluctuations caused by a fluctuation in $\theta$ at fixed $\varphi$. Below we show that keeping $\delta$ small ensures the agreement of the Glauber and Langevin steady state.
In the steady state, the phase-space distribution is stationary. Hence, the probability fluxes into and out of a given state balance out:
$\left[\rho(\theta_{i}+\delta,\varphi_{i}+\delta)w(\theta_{i}+\delta\rightarrow\theta_{i})-\rho(\theta_{i},\varphi_{i})w(\theta_{i}\rightarrow\theta_{i}+\delta)\right]\Big{|}_{\text{constr.}}=0\ ,$ (90)
where the specific form of the Glauber transition rate at constant $\varphi_{i}$ is

### v4

@@ D.1 Langevin steady-state distribution
with conjugate degrees of freedom $\{\theta_{i},\varphi_{j}\}{=}\delta_{ij}$. To maintain consistency with the traditional form of Langevin dynamics and represent the equations of motion on the constraint surface, the notation $\partial_{\theta_{i}}H|_{\text{constr.}}$ is introduced to replace the Poisson bracket. Here $\eta_{i}(t)$ denotes white noise with $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$, with $T$ the temperature of the bath, where the average is taken over the phase-space density $\rho(\vec{\theta}(t),\vec{\varphi}(t))$. Note that the constraint implies that each pair $(\theta,\varphi_{i})$ experiences the same noise realization.
Below, we first consider steady-state distributions (i.e., no explicit $t$-dependence in $\rho$, Sec. D.1.1), and then generalize the result to non-stationary distributions in Sec. D.1.2.
D.1.1 Steady-state distribution
To obtain the steady-state phase-space density distribution $\rho(\vec{\theta}(t),\vec{\varphi}(t))$ under the above dynamics, we follow the standard procedure [65, 72] and introduce a test function $f(\vec{\theta}(t),\vec{\varphi}(t))$ over phase space. The change of the test function within a small time window $\Delta t$ is to leading order
$f\left(\vec{\theta}(t+\Delta t),\vec{\varphi}(t+\Delta t)\right)-f\left(\vec{\theta}(t),\vec{\varphi}(t)\right)$ (146)
@@ D.1.1 Steady-state distribution
$\int\frac{\mathrm{d}f(\vec{\theta},\vec{\varphi})}{\mathrm{d}t}\rho(\vec{\theta},\vec{\varphi})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=\lim_{\Delta t\to 0}\int\frac{f\left(\vec{\theta}(t+\Delta t),\vec{\varphi}(t+\Delta t)\right)-f\left(\vec{\theta}(t),\vec{\varphi}(t)\right)}{\Delta t}\rho(\vec{\theta},\vec{\varphi})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=0.$ (148)
To proceed with the derivation, we write the noise terms in Eq. (D.1) as a Wiener process $W_{i}(t)$:
@@
$\int\frac{\partial f}{\partial\theta_{i}}\dot{\theta_{i}}\Delta t\;\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}\mathbf{\theta}_{i}\mathrm{d}\mathbf{\varphi}_{i}$ $=$ $\int\frac{\partial f}{\partial\theta_{i}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$ (150)
$=$ $\Delta t\int\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}},$
@@
$\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\dot{\theta_{i}}\dot{\theta_{j}}\Delta t^{2}\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}=$ (151)
$=$ $\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\left[\frac{\partial H}{\partial\varphi_{j}}\Bigg|_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{j}(t)\right]\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\int 2T\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\Delta W_{i}(t)\Delta W_{j}(t)\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{j}}\Bigg|_{\text{constr.}}\Delta W_{i}(t)\rho\left(\vec{\theta},\vec{\varphi}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\Delta W_{j}(t)\rho\left(\vec{\theta},\vec{\varphi}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}+\Delta t^{2}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\frac{\partial H}{\partial\varphi_{j}}\Bigg|_{\text{constr.}}\right)\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\Delta t\delta_{ij}2T\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\rho\left(\vec{\theta},\vec{\varphi}\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+O\left[\left(\Delta t\right)^{2}\right],$
@@
$+T\left(\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{i}}+\frac{\partial^{2}f}{\partial\varphi_{i}\partial\varphi_{i}}+2\frac{\partial^{2}f}{\partial\varphi_{i}\partial\theta_{i}}\right)\Bigg]\rho\left(\vec{\theta},\vec{\varphi}\right)$ $=$ $0\,.$ (152)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[\frac{\partial f}{\partial\theta_{i}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg|_{\text{constr.}}\right)-\frac{\partial f}{\partial\varphi_{i}}\left(\frac{\partial H}{\partial\theta_{i}}\Bigg|_{\text{constr.}}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\vec{\theta},\vec{\varphi}\right)=0.$ (153)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[-\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)f\frac{\partial H}{\partial\theta_{i}}\Bigg|_{\text{constr.}}+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\vec{\theta},\vec{\varphi}\right)=0.$ (154)
@@
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}f\left[\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left(\frac{\partial H}{\partial\theta_{i}}\Bigg|_{\text{constr.}}\rho\left(\vec{\theta},\vec{\varphi}\right)\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}\rho\left(\vec{\theta},\vec{\varphi}\right)\right]=0.$ (155)
@@
$\sum_{i}\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left[\frac{\partial H}{\partial\theta_{i}}\Bigg|_{\text{constr.}}\ \rho\left(\vec{\theta},\vec{\varphi}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\rho\left(\vec{\theta},\vec{\varphi}\right)\right]=0.$ (156)
So far the derivation was generic since we did not make use of the precise form of the equations of motion. We will show in Sec. D.2 that Eq.  (156) governs also the constrained Glauber dynamics.
D.1.2 Time-dependent non-stationary distribution
Before we do that, let us complete the analysis and examine the case where the density distribution exhibits explicit time dependence, $\rho(\vec{\theta}(t),\vec{\varphi}(t),t)$. Such situations can occur away from equilibrium where oscillating states arise.
@@ D.1.2 Time-dependent non-stationary distribution
Setting Eq.(D.1.2) equal to Eq.(159) yields
@@
This Fokker-Planck equation, consistent with Eq. (144), governs the dynamics of the phase-space distribution. As expected, for a stationary distribution, $\partial_{t}\rho=0$, it reduces to Eq. (156).
D.1.3 Eliminating the auxiliary degree of freedom from Langevin dynamics
When studying the nonreciprocal Langevin dynamics, it is often convenient to keep only the original angles $\vec{\theta}$ and not track the auxiliary variables $\vec{\varphi}$ explicitly. In this reduced description, the dynamics is written directly in terms of the (generally nonreciprocal) drift force $F_{i}(\vec{\theta})$ acting on $\theta_{i}$.
We start from the overdamped Langevin equation (Itô convention)
$\theta_{i}(t+\Delta t)-\theta_{i}(t)=F_{i}(\vec{\theta})\,\Delta t+\sqrt{2T}\,\Delta W_{i}(t),$ (162)
where $\Delta W_{i}$ are Wiener increments with $\langle\Delta W_{i}\rangle=0$ and $\langle\Delta W_{i}\,\Delta W_{j}\rangle=\delta_{ij}\Delta t$. The corresponding configuration-space probability density $\rho(\theta_{i},t)$ obeys the standard Fokker–Planck equation
$\partial_{t}\rho(\vec{\theta},t)=-\sum_{i}\partial_{\theta_{i}}\left[F_{i}(\vec{\theta})\,\rho(\vec{\theta},t)\right]+T\sum_{i}\partial_{\theta_{i}}^{2}\rho(\vec{\theta},t).$ (163)
In a stationary state, $\partial_{t}\rho=0$, and we obtain
$0=\sum_{i}\partial_{\theta_{i}}\left[F_{i}(\vec{\theta})\,\rho(\vec{\theta})+T\,\partial_{\theta_{i}}\rho(\vec{\theta})\right].$ (164)
For the vision-cone XY model in Eq. (141), the drift force is
$F_{i}(\vec{\theta})=-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\,\sin(\theta_{i}-\theta_{j}).$ (165)
This drift force is precisely the one obtained from the Hamiltonian embedding once the constraint $\varphi_{i}=\theta_{i}+\pi$ is imposed:
$F_{i}(\vec{\theta})=\partial_{\varphi_{i}}H\Big|_{\mathrm{constr.}}=-\partial_{\theta_{i}}H\Big|_{\mathrm{constr.}}.$ (166)
Moreover, the higher-order derivatives of $F_{i}(\vec{\theta})$ with respect to $\theta_{i}$ can be expressed directly in terms of the constrained Hamiltonian:
$\partial_{\theta_{i}}^{n}F_{i}(\vec{\theta})=-\partial_{\theta_{i}}\big(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\big)^{n}H\Big|_{\mathrm{constr.}},\qquad n=1,2,\ldots.$ (167)
Thus, we can always describe the force $F_{i}$ and its higher-order derivatives with respect to $\theta_{i}$ using the Hamiltonian embedding. The steady-state distribution and non-stationary dynamics of the density introduced in Eq. (156) and Eq. (161) is recovered by expressing all derivatives of $\rho(\vec{\theta})$ in Eq. (163) and Eq. (164), $\partial_{\theta_{i}}^{n}\rho(\vec{\theta})$, in terms of $(\partial_{\theta_{i}}+\partial_{\varphi_{i}})^{n}\rho(\vec{\theta},\vec{\varphi})$. This shows that the Langevin dynamics obtained from Hamiltonian embedding on the constrained manifold and the conventional nonreciprocal Langevin description give the same evolution for the physical degrees of freedom.
D.2 Glauber dynamics description
We now present a Glauber algorithm that samples the steady state and the nonstationary distribution of the Langevin dynamics using the Hamiltonian embedding. In particular, we focus on defining transition rates that are consistent with the constraint.
When the Hamiltonian is evaluated on the constraint, it vanishes identically, $H(\vec{\theta},\vec{\varphi})\big|_{\text{constr.}}=0$. Thus, it is nontrivial to define the Glauber update rule based on energy differences alone. To address this, we consider a Markov process in which every state $\{\vec{\theta},\vec{\varphi}\}$ satisfies the constraint $\theta_{i}=\varphi_{i}+\pi$ while the transition rates are defined based on the energy difference between two states that differ by a small increment $\delta$ in the angle of the original spin and the auxiliary spin, respectively.
We show that taking $\delta$ sufficiently small guarantees that the Glauber and Langevin steady states coincide. We emphasize that, in the strict limit $\delta\to 0$, the Hamiltonian vanishes identically under the constraint, making the update rule ill-defined. However, for any small but finite $\delta$, the Glauber dynamics correctly reproduces the physics of the nonreciprocal Langevin equation.
We start from the probability current from state $\{\vec{\theta},\vec{\varphi}\}$ to state $\{\theta_{i}+\delta,\{\vec{\theta}\}_{j\neq i},\varphi_{i}+\delta,\{\vec{\varphi}\}_{j\neq i}\}$, which is given by
$\mathcal{J}(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)$ $=\rho(\vec{\theta},\vec{\varphi})w(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)$ (168)
$-\rho(\theta_{i}+\delta,\{\vec{\theta}\}_{j\neq i},\varphi_{i}+\delta,\{\vec{\varphi}\}_{j\neq i})w(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i}).$
Here the increment $\delta$ is restricted to the interval $\Delta,\Delta]$. The Glauber transition rates read as